OpenStudy (anonymous):

What are the 3 resonce CNO- least to most important?

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@BAdhi

OpenStudy (anonymous):

@Koikkara

OpenStudy (koikkara):

hmm, i think -_ + - .C=N-O. - ..C=N-O - ..C=N=O

OpenStudy (koikkara):

here, all of them have at least one reactive electron on carbon and the bottom two have two reactive electrons on carbon. I believe that the top structure is the most stable although the separation in formal charges

OpenStudy (koikkara):

ref : https://www.youtube.com/watch?v=HxU53Ignz5k

OpenStudy (anonymous):

Thanks. Would you like for me draw it out to double check it?

OpenStudy (koikkara):

Check this pdf.

OpenStudy (anonymous):

http://oi61.tinypic.com/rw2hd5.jpg

OpenStudy (anonymous):

That's how I did them

OpenStudy (anonymous):

@Shalante

OpenStudy (anonymous):

@Hero